OpenStudy (anonymous):

Which expression is equivalent to 2^4 ⋅ 2^−7? 1 over 2 to the power of 11 1 over 2 to the power of 3 2^3 2^11

OpenStudy (anonymous):

@Frozen15

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (frozen15):

what were you thinking as an answer or guess?

OpenStudy (xapproachesinfinity):

i figured that's what you meant

OpenStudy (anonymous):

i dont know how to do 2 to the power of negative seven

OpenStudy (xapproachesinfinity):

\(2^4.2^{-7}=2^{4-7}\)

OpenStudy (xapproachesinfinity):

the property been used is \(a^na^m=a^{a+m}\)

OpenStudy (xapproachesinfinity):

simplify that

OpenStudy (anonymous):

i understand that

OpenStudy (anonymous):

but how do i do it to the power of a negative

OpenStudy (xapproachesinfinity):

ok tell me what diid you get for that first

OpenStudy (anonymous):

2 to the power of 4 ?

OpenStudy (xapproachesinfinity):

no how come? 4-7=-3 so it will be 2^-3

OpenStudy (anonymous):

negative seven

OpenStudy (xapproachesinfinity):

you forgot how to add numbers

OpenStudy (anonymous):

4 - (-7) equals -3

OpenStudy (anonymous):

i dont understand how

OpenStudy (xapproachesinfinity):

see what is says there 2 raised to 4-7 that means do the addition of 4-7 to get -3 so 2^-3

OpenStudy (anonymous):

4 - (-7) equals 4 + 7

OpenStudy (anonymous):

so 11

OpenStudy (xapproachesinfinity):

now you got \(2^{-3}\) you can use this rule \(a^{-n}=1/a^n\)

OpenStudy (anonymous):

it cant be negative three

OpenStudy (anonymous):

how do u have a smart score of 94 if you don-t know this

OpenStudy (xapproachesinfinity):

no where did you get that other negative in the middle don 't just write things up from your head we are given 4-7 which equals -3

OpenStudy (xapproachesinfinity):

because i'm dumb, if i know i would be ignorant

OpenStudy (anonymous):

i said 2 to the power of negative seven

OpenStudy (xapproachesinfinity):

yes and that gives 2^4 times 2^-7 equals 2^ (4-7)

OpenStudy (xapproachesinfinity):

the rule is described above

OpenStudy (anonymous):

i don't have that as an option

OpenStudy (anonymous):

i am doing multiple choice

OpenStudy (anonymous):

it has to be 11

OpenStudy (xapproachesinfinity):

we are getting there

OpenStudy (anonymous):

idk how

OpenStudy (xapproachesinfinity):

i was trying to explain but you accused me of not been capalble of my smart score

OpenStudy (anonymous):

how did u get negative three from 4 and negative seven

OpenStudy (xapproachesinfinity):

first we obtain 2^-3 from 2^(4-7) now 2^-3 =1/2^3

OpenStudy (anonymous):

you lost you marbles

OpenStudy (xapproachesinfinity):

you owe me 4 dollar and i owe you 7 the result is that i still owe you 3

OpenStudy (anonymous):

right now your head is a piping jalapeno

OpenStudy (xapproachesinfinity):

that explain why 4-7=-3

OpenStudy (anonymous):

negative 7

OpenStudy (anonymous):

4 minus negative seven

OpenStudy (anonymous):

4 - (-7)

OpenStudy (anonymous):

what don't u understand by negative seven

OpenStudy (anonymous):

is it in your vocabulary

OpenStudy (anonymous):

cmon